# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

### v2

@@ 7.2.3 Perceptual Study
Figure 9: THUNDER-F perceptual study - dynamism The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER THUNDER-F w/o mesh-to-speech. However, ground truth is still rated more dynamic.
@@
Figure 10: THUNDER-F perceptual study - realism The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER-F w/o mesh-to-speech. THUNDER-T w/o mesh-to-speech is rated about as realistic as THUNDER. Remarkably, the participants have preferred THUNDER’s generation over ground truth in terms of realism..
@@
Figure 11: THUNDER-T complete perceptual study. The results of THUNDER-T against other methods. THUNDER-T outperforms its counterpart with mesh-to-speech and the deterministic FlameFormer on lip-sync quality. Additionally, the participatns have a slight preference for THUNDER-T over ground truth. This suggests, that while THUNDER-T exhibits a degree of dynamis, the GT potential has not yet been fully exhausted.
@@
Figure 12: THUNDER-T perceptual study - dynamism. The participants rate THUNDER-T to be more dynamic than THUNDER-T without mesh-to-speech and also more dynamic than FlameFormer-T, THUNDER-T w/o mesh-to-speech and even Ground Truth. This suggest that THUNDER-T’s realism exhausted the realism of the dataset.
@@
Figure 13: THUNDER-T perceptual study - realism. The participants prefer the degree of realism of THUNDER-T’s outputs over FlameFormer-T
@@
Here we provide the rest of our comprehensive perceptual evaluation of THUNDER. We run a perceptual study on Amazon Mechanical Turk. The participants are shown two videos side-by-side. One corresponding to THUNDER and one to a baseline. The left-right order is randomized. The participants must finish watching both videos before being allowed to rate the videos on a five-point Likert scale (strong preference for left, weak preference for left, indifference, weak preference for right and strong preference for right). We ask the participants to rate three aspects of the animation - lip-sync, dynamism and realism. The exact task description can be seen in the study template in Fig. 14. The participants are shown 20 comparisons, generated from test audios, which where randomly selected from the RAVDESS and TCD-TIMIT test sets. In addition to that, we repeat the first 4 comparisons at the end of the study and discard the first 4 responses. This gives the participants a few examples such that they can adjust to the task. Additionally, we include 4 catch trials, where one animation is selected from Ground Truth and the other animation is clearly wrong. We discard the participants that are wrong on more than one catch trial. The complete results of the perceptual study for THUNDER for lip-sync, dynamism and realism can be found in Figures 10, 10 and 10 respectively. The results for THUNDER-T are given in Figures 13, 13 and 13. The results validate our qualitative and quantiative findings - THUNDER models exhibit better lip-sync, realism and dynamism when compared to baselines.
@@
7.2.4 Generalization to other datasets
To verify whether the performance of THUNDER and mesh-to-speech translates to other types of data, we run additional experimetns on TFHP (the DiffPoseTalk dataset [sun2024diffposetalk]. TFHP was reconstructed from YouTube videos The dataset contains unscripted audios, longer sequences, and the reconstructions were obtained using MICA [MICA:ECCV2022] and SPECTRE [filntisis2022visual] and look qualitatively very different compared to our EMICA reconstructions. Can THUNDER work also in this setting?
We train both mesh-to-speech and THUNDER-T on TFHP, along with training DiffPoseTalk with the official DiffPoseTalk codebase. Please note, that in order to equalize the input conditions, we trained DiffPoseTalk without identity condioning. During evaluation, we do not use the style-condition but evaluate using the audio condition only (for both THUNDER models and DiffPoseTalk). The results are reported in Tab. 6.
Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ T-FDD-U $\downarrow$
DiffPoseTalk (w/o style cond.) 1.24 0.341 0.508 0.25 0.0388 0.0395 0.235 0.318 0.0111
THUNDER-T w/o m2s 1.1 0.332 0.518 0.226 0.0208 0.0293 0.122 0.277 0.0068
THUNDER-T 1.09 0.351 0.545 0.214 0.0255 0.0322 0.103 0.265 0.00777
@@ 7.2.4 Generalization to other datasets
Comparing different speech-driven animation methods is a difficult task, especially if trained on a different dataset and especially if the data was acquired in different ways (4D scans, pseudo-GT of different reconstruction methods, etc.). Having a fair apples-to-apples comparisons stemming from measurements done against the (pseudo-) ground truth is therefore difficult. Regardless, for the sake of completeness, we compare THUNDER to the official releases of recent methods that were also trained on pseudo-GT, namely EMOTE [danecek2023emote] (a deterministic SOTA) and DiffPoseTalk [sun2024diffposetalk] (stochastic SOTA). EMOTE is a FaceFormer-inspired architecture with additional input conditions for emotions and intensity, trained with a content-emotion disentanglement mechanism. It was trained on pseudo-GT reconstructions of MEAD. The pGT was acquired with the same INFERNO tracker, making the resulting metrics comparable. The macro-architecture of DiffPoseTalk’s diffusion is very similar to THUNDER and apart from the additional style conditioning input, it differs only in small details (such as noise schedule, number of denoising steps, etc.). The most important difference is that it was trained on a different dataset, and the reconstructions were produced by a combination of SPECTRE [filntisis2022visual] and MICA [MICA:ECCV2022] and hence look qualitatively very different. As such, the DiffPoseTalk metrics likely dominated by the qualitative difference, and are only listed for completeness. Please note that for this experiment, DiffPoseTalk was run without the style condition (to match the setting of THUNDER). Finally, we include a FlameFormer trained on our data, and THUNDER trained without mesh-to-speech. The results are reported in Tab. 7.
Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ FDD-U $\downarrow$ FDD-L $\downarrow$
FlameFormer* 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
EMOTE 1.06 0.221 0.442 0.466 0.0563 0.0401 0.248 0.24 0.0141 0.104
DiffPoseTalk 1.15 0.255 0.38 0.316 0.0345 0.0479 0.226 0.308 0.0141 0.102
THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
THUNDER 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
@@ 7.2.6 Qualitative Results
Diversity Input speech may originate from many different expression and even emotions. A well-trained stochastic model should be able to account for that and generate multiple plausible animations, possibly with different facial expressions. Fig. 15 demonstrates THUNDER’s ability to do so.
@@ 8 Architecture of Baselines
Figure 16: Architecture of re-implemented baselines. (a) Our version of Media2Face [zhao2024media2face] utilized the same architecture as THUNDER but takes an extra ImabeBind feature on the input. (c) Our re-implementation of DiffPoseTalk [sun2024diffposetalk] is also based on THUNDER and takes an extra style feature from a pretrained style encoder. (c) FlameFormer is an adapted FaceFormer [fan2022faceformer]. FlameFormer predicts FLAME expression instead of full vertex space. Furthermore, FlameFormer adapts a non-autoregressive BERT-like prediction mechanism (akin to EMOTE [danecek2023emote]).
@@
Media2Face. Our reimplementation of Media2Face is built on our diffusion architecture. In training, we provide additional input of ImageBind[imagebind_Girdhar2023] features extracted from images of the corresponding videos of the training set. Like the audio condition, this condition is dropped 20% of the time.
DiffPoseTalk. Our reimplementation of DiffPoseTalk is also built on our diffusion architecture. It is trained in two stages. In the first stage we train the contrastive style encoder as proposed by the authors [sun2024diffposetalk]. In training of DiffPoseTalk, we provide the additional style feature on the input. The style feature is extracted from the GT animation sequence. Like the audio condition, the style condition is dropped 20% of the time.
FlameFormer. Our FlameFormer architecture is based on FaceFormer. Like FaceFormer, FlameFormer is a transformer-based deterministic speech-driven animation network that takes addition categorical style condition on the input. Instead of predicting the full vertex space like the original FaceFormer does, FlameFormer predicts FLAME 3DMM coefficients $\boldsymbol{x}$. Additionally, instead of of using the autoregressive loop with a transformer decoder, FlameFormer’s decoder is a transformer encoder, which eliminates the need for the autoregressive formulation. This does not degrade performance and makes the model more efficient, as was already shown by Daněček et al. [danecek2023emote], which have proposed the FlameFormer baseline in their paper. Furthermore, we extend FlameFormer’s conditioning inputs - the subject identity (one-hot vector) is complemented with a one-hot vector for 8 basic emotions and three intensities.
@@ 9.1 Choice of methodology
Acquiring high quality 4D scans of sufficient scale and richness is costly, time-consuming and requires specialized hardware. Hence, recent speech-driven animation works [sun2024diffposetalk, danecek2023emote, zhao2024media2face, yang2024probabilistic] have turned to pseudo-GT recovered from videos. This choice comes with an important decision - which face reconstruction methodolgy to use. There are several options but two of the most viable options are:
@@
1) Optimization-based 3DDM fitting with analysis-by-image-synthesis [VetterBlanz1998, Blanz2002, Blanz2003]. While this options is viable and would produce good results, this optimization-based process is rather inefficient when deployed to many hours of video.
@@
2) Off-the-shelf face 3DMM regressors. Due to the computational intensity of optimization-based fitting and thanks to the recent advance in deep-learning based in-the-wild face reconstruction methods [Sanyal2019_RingNet, Feng2021_DECA, EMOCA:CVPR:2021, MICA:ECCV2022, filntisis2022visual, zhang2023tokenface, Retsinas_2024_CVPR_Smirk, Tewari2017, Tewari2018, Tewari2019_FML, Deng2019]. While this is still an active area of research, and no method produces results that could be considered as accurate as 4D scans, recent years have brought enough advancement so that they can be employed for speech-driven animation dataset construction.
9.2 Discussion of FLAME regressors
@@
Since THUNDER utilizes FLAME [FLAME:SiggraphAsia2017], our discussion will focus on FLAME regressors (trackers) only. There are several systems that have been used for talking head avatar research. No tracker Each tracker produces different kinds of errors but they were used in prior works on 3D talking head avatars. Here we briefly discuss some of the most applicable candidates:
  • DECA [Feng2021_DECA] is the first FLAME-based regressor trained with the self-supervised ”analysis-by-image-synthesis” loop. However, it is limited in terms of richness of expressions and quality of lip animations.
@@
  • EMOCA [EMOCA:CVPR:2021] is a follow-up of DECA, capable of reconstruct rich emotions thanks to its emotion-consistency loss. However it can exaggarate expressions and the lip-sync is not good enough.
@@
  • SPECTRE [filntisis2022visual] applies a paradigm similar to EMOCA but leverages a lip-reading consistency loss instead. It produces good lip-readable animations, but unfortunately often exaggarates, is not capable of reconstructing rich emotional expressions and the disentanglement between expression and identity is rather poor.
@@
  • EMICA [inferno2023] is a combination of the DECA, EMOCA and SPECTRE putting together the benefits of them all to achieve high quality emotions and lip-sync. Additionally, EMICA makes use of MICA [MICA:ECCV2022] for a consistent identity prediction. The system is trained jointly and publicly available.
@@
  • SMIRK [Retsinas_2024_CVPR_Smirk] SMIRK is a recent reconstruction method which uses a neural renderer and in-the-loop neural data augmentation to produce high quality reconstructions.
9.3 Previous works
@@
As a result of the tremendous progress of in-the-wild face reconstruction, some of the recent SOTA speech-driven animation have already turned to pseudo-GT as the primary source of data.
  • EMOTE [danecek2023emote] employed the EMICA model to reconstruct the MEAD [kaisiyuan2020mead] dataset.
  • Yang et al. [yang2024probabilistic] utilized both DECA and SPECTRE to reconstruct LRS3 [afouras2018lrs3].
@@
  • DiffPoseTalk employed MICA and SPECTRE separately and then combined the shape predictions of MICA and expression predictions of SPECTRE to reconstruct TFHP.
  • EMOCA exaggarates expressions.
  • DECA’s lip-sync and expressions are very limited.
@@
  • We opt for EMICA as it provides a good compromise - rich emotions, good lip animations, consistent shape identity and temporal consistency. That said, it is not artifact-free (mouth wide open correlates with raised eyebrows, etc.) but it is sufficient to use for our research.
Following EMOTE, we opt to use EMICA in this work as well.

### v3

@@ 7.2.4 Perceptual Study
Figure 8: THUNDER-F perceptual study - dynamism. The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER-F w/o mesh-to-speech. However, ground truth is still rated more dynamic than THUNDER-F.
@@
Figure 9: THUNDER-F perceptual study - realism. The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER-F w/o mesh-to-speech. THUNDER-T w/o mesh-to-speech is rated about as realistic as THUNDER. Remarkably, the participants have preferred THUNDER’s generation over ground truth in terms of realism, which suggests the training dataset has been exhausted.
@@
Figure 10: THUNDER-T perceptual study - lip-sync. The results of THUNDER-T against other methods. THUNDER-T outperforms its counterpart with mesh-to-speech and the deterministic FlameFormer on lip-sync quality. Additionally, the participants have a slight preference for THUNDER-T over ground truth. This suggests, that THUNDER-T saturates the lip-sync quality of our pseudo-GT.
@@
Figure 11: THUNDER-T perceptual study - dynamism. The participants rate THUNDER-T to be more dynamic than THUNDER-T without mesh-to-speech and also more dynamic than FlameFormer-T and THUNDER-T w/o mesh-to-speech. Pseudo-GT, however, is preferred over THUNDER-T, suggesting that THUNDER-T’s dynamism did not yet exhaust the dynamism of the training dataset.
@@
Figure 12: THUNDER-T perceptual study - realism. The participants prefer the degree of realism of THUNDER-T’s outputs over the deterministic FlameFormer-T and diffusion-based THUNDER without mesh-to-speech.
@@
Here we provide the rest of our comprehensive perceptual evaluation of THUNDER. We run a perceptual study on Amazon Mechanical Turk. The participants are shown two videos side-by-side. One corresponding to THUNDER and one to a baseline. The left-right order is randomized. The participants must finish watching both videos before being allowed to rate the videos on a five-point Likert scale (strong preference for left, weak preference for left, indifference, weak preference for right and strong preference for right). We ask the participants to rate three aspects of the animation - lip-sync, dynamism and realism. The exact task description can be seen in the study template in Fig. 13. The participants are shown 20 comparisons, generated from test audios, which where randomly selected from the RAVDESS and TCD-TIMIT test sets. In addition to that, we repeat the first 4 comparisons at the end of the study and discard the first 4 responses. This gives the participants a few examples such that they can adjust to the task. Additionally, we include 4 catch trials, where one animation is selected from Ground Truth and the other animation is clearly wrong. We discard the participants that are wrong on more than one catch trial. The complete results of the perceptual study for THUNDER for lip-sync, dynamism and realism can be found in Figures 9, 9 and 9 respectively. The results for THUNDER-T are given in Figures 12, 12 and 12. The results validate our qualitative and quantiative findings - THUNDER models exhibit better lip-sync, realism and dynamism when compared to baselines.
@@
Comparing different speech-driven animation methods is a difficult task, especially if trained on a different dataset and especially if the data was acquired in different ways (4D scans, pseudo-GT of different reconstruction methods, etc.). Having a fair apples-to-apples comparisons stemming from measurements done against the (pseudo-) ground truth is therefore difficult. Regardless, for the sake of completeness, we compare THUNDER to the official releases of recent methods that were also trained on pseudo-GT, namely EMOTE [18] (a deterministic SOTA) and DiffPoseTalk [74] (stochastic SOTA). EMOTE is a FaceFormer-inspired architecture with additional input conditions for emotions and intensity, trained with a content-emotion disentanglement mechanism. It was trained on pseudo-GT reconstructions of MEAD. The pGT was acquired with the same INFERNO tracker, making the resulting metrics comparable. The macro-architecture of DiffPoseTalk’s diffusion is very similar to THUNDER and apart from the additional style conditioning input, it differs only in small details (such as conditioning with the FLAME shape vector, noise schedule, number of denoising steps, etc.). The most important difference is that it was trained on a different dataset, and the reconstructions were produced by a combination of SPECTRE [29] and MICA [100] and hence look qualitatively very different. As such, the DiffPoseTalk metrics likely dominated by the qualitative difference, and are only listed for completeness. Please note that for this experiment, DiffPoseTalk was run without the style condition (to match the setting of THUNDER). Finally, we include a FlameFormer trained on our data, and THUNDER trained without M2S. The results are reported in Tab. 8.
Lip-Sync Sample diversity Temporal diversity Face Dynamic Dev.
Name Dataset Input LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ FDD-U $\downarrow$ FDD-L $\downarrow$
FlameFormer* THUNDERSET audio, one-hot speaker ID 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
EMOTE MEAD audio, one-hot speaker ID, emotion and intensity 1.06 0.221 0.442 0.466 0.0563 0.0401 0.248 0.24 0.0141 0.104
DiffPoseTalk TFHP audio, shape vector 1.15 0.255 0.38 0.316 0.0345 0.0479 0.226 0.308 0.0141 0.102
THUNDER-F w/o m2s THUNDERSET audio only 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
THUNDER-F THUNDERSET audio only 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
@@ 7.2.6 Qualitative Results
In this section we provide additional qualitative results and comparisons. Furthermore, we strongly recommend watching the supplemental video for an audible and in-motion qualitative results and comparisons.
Qualitative Comparison on THUNDERSET. Fig. 14 is a super-set of the figure listed in the main paper (Fig. 5).
Figure 14: Qualitative comparison on THUNDERSET. This figure shows the comparison between baselines, our model and GT for selected utterances. Note that DiffPoseTalk was trained on TFHP and EMOTE on MEAD. Supplemental PDF and video contain more qualitative comparisons.
Diversity. Input speech may originate from many different expression and even emotions. A well-trained stochastic model should be able to account for that and generate multiple plausible animations, possibly with different facial expressions. Fig. 15 demonstrates THUNDER’s ability to do so.
@@
Qualitative Comparison on TFHP. Here we provide qualitative results of models trained on TFHP (the same models from Sec. 7.2.1). We select a few utterances from various TFHP test-set sequences. We show results of models trained without head pose in Fig. 16, the unposed results of models trained with pose in Fig. 17, and finally the same results with the generated pose in Fig. 18.
Figure 16: TFHP models trained without head pose. We
Figure 17: Results of TFHP-trained models, trained with head pose. The results are shown in canonical space, i. e.  without the generated head pose.
Figure 18: TFHP models trained with head pose. The results are shown posed with the predicted pose (or pseudo-GT pose in case of pseudo-GT).
@@ 8 Architecture of Baselines
Figure 19: Architecture of re-implemented baselines. (a) Our version of Media2Face [98] utilized the same architecture as THUNDER but takes an extra ImabeBind feature on the input. (c) Our re-implementation of DiffPoseTalk [74] is also based on THUNDER and takes an extra style feature from a pretrained style encoder. (c) FlameFormer is an adapted FaceFormer [26]. FlameFormer predicts FLAME expression instead of full vertex space. Furthermore, FlameFormer adapts a non-autoregressive BERT-like prediction mechanism (akin to EMOTE [18]).
@@
Media2Face. Our reimplementation of Media2Face is built on our diffusion architecture. In training, we provide additional input of ImageBind [32] features extracted from images of the corresponding videos of the training set. Like the audio condition, this condition is dropped 20% of the time.
DiffPoseTalk. Our reimplementation of DiffPoseTalk is also built on our diffusion architecture. It is trained in two stages. In the first stage we train the contrastive style encoder as proposed by the authors [74]. In training of DiffPoseTalk, we provide the additional style feature on the input. The style feature is extracted from the GT animation sequence. Like the audio condition, the style condition is dropped 20% of the time.
FlameFormer. Our FlameFormer architecture is based on FaceFormer. Like FaceFormer, FlameFormer is a transformer-based deterministic speech-driven animation network that takes addition categorical style condition on the input. Instead of predicting the full vertex space like the original FaceFormer does, FlameFormer predicts FLAME 3DMM coefficients $\boldsymbol{x}$. Additionally, instead of of using the autoregressive loop with a transformer decoder, FlameFormer’s decoder is a transformer encoder, which eliminates the need for the autoregressive formulation. This does not degrade performance and makes the model more efficient, as was already shown by Daněček et al. [18], which have proposed the FlameFormer baseline in their paper. Furthermore, we extend FlameFormer’s conditioning inputs - the subject identity (one-hot vector) is complemented with a one-hot vector for 8 basic emotions and three intensities.
@@ 9.1 Choice of methodology
Acquiring high quality 4D scans of sufficient scale and richness is costly, time-consuming and requires specialized hardware. Hence, recent speech-driven animation works [74, 18, 98, 95] have turned to pseudo-GT recovered from videos. This choice comes with an important decision - which face reconstruction methodolgy to use. There are several options but two of the most viable options are:
@@
1) Optimization-based 3DDM fitting. Approaches like these (for instance fitting with analysis-by-image-synthesis) have been proposed more than two decades ago [87, 5, 6]. While these options are viable and would produce good results, the optimization-based process is rather inefficient when deployed to many hours of video. This remains true even for the contemporary optimization-based trackers, such as the MICA Metrical Tracker [100] or Pixel3DMM [31].
@@
2) Off-the-shelf face 3DMM regressors. Due to the computational intensity of optimization-based fitting and thanks to the recent advance in deep-learning based in-the-wild face reconstruction methods [72, 28, 17, 100, 29, 97, 68, 80, 81, 82, 21]. While this is still an active area of research, and no method produces results that could be considered as accurate as 4D scans, recent years have brought enough advancement so that they can be employed for speech-driven animation dataset construction.
9.2 Discussion of existing FLAME regressors
@@
Since THUNDER utilizes FLAME [51], our discussion will focus on FLAME regressors (trackers) only. There are several systems that have been used for talking head avatar research. No tracker is perfect. Each tracker produces different kinds of errors but despite that, the recent trackers are finally good enough to be used for talking head avatar research. In fact, a few previous methods have already made use of them [18, 55, 74]. Here we briefly discuss some of the most applicable candidates:
DECA [28] is the first FLAME-based regressor trained with the self-supervised ”analysis-by-image-synthesis” loop. While this paper was SOTA at the time of its release, it still had considerable artifacts, especially when it comes to the richness of facial expression and emotions and quality of lip animations. Furthermore, the stability of the identity prediction was not stable when applied to videos.
@@
EMOCA [17] is a follow-up of DECA, capable of reconstructing rich emotions thanks to its emotion-consistency loss. However, it can exaggarate expressions and the lip-sync is not good enough. EMOCAv2 later addressed the lip-sync issue by incorporating the perceptual lip reading loss from SPECTRE, obtaining the best of both worlds. However, similarly to DECA and SPECTRE, EMOCAv2 exhibits unstable identity prediction over the course of a video.
@@
SPECTRE [29] applies a paradigm similar to EMOCA but leverages a lip-reading consistency loss instead. It produces good lip-readable animations, but unfortunately often exaggarates and over-articulates, is not capable of reconstructing rich emotional expressions and the disentanglement between expression and identity is rather poor, which often makes the predicted expression vector compensate the inacuraccies in identity predictions.
@@
EMICA [19] is a combination of several methods’ contributions. It combines the benefits of DECA, EMOCA and SPECTRE to achieve high quality emotions and lip-sync. Additionally, EMICA makes use of MICA [100] for a consistent identity prediction. The system is trained jointly and publicly available.
@@
SMIRK [68] SMIRK is a recent reconstruction method which uses a neural renderer and in-the-loop neural data augmentation to produce high quality reconstructions.
9.3 Data acquisition in previous works
@@
As a result of the tremendous progress of in-the-wild face reconstruction, some of the recent SOTA speech-driven animation methods have already turned to pseudo-GT as the primary source of data. For instance:
EMOTE [18] employed the EMICA model to reconstruct the MEAD [89] dataset.
Yang et al. [95] utilized both DECA and SPECTRE to reconstruct LRS3 [1].
@@
DiffPoseTalk [74] employed MICA and SPECTRE separately and then combined the shape predictions of MICA and expression predictions of SPECTRE to reconstruct the TFHP dataset.
@@
We opt for EMICA as it provides a good compromise - rich emotions, good lip animations, consistent shape identity and temporal consistency of the identity prediction. While not artifact-free, EMICA is a good option for building a large-scale dataset for 3D talking head avatar research.